22 January, 2017

19:00

Observatory was closed.. Roshaan and i had a good conversation on Astronomy, life and everything in between. 

21 January, 2017

20:00

Cloudy and humid weather is keeping the observatory room closed.  It will continue for a few more days.

I am trying the USB coonection from the observatory to the meeting room. Cat5/6 cable connection is not working. Have checked two different cables and connectors. This is a problem that needa to be resolved ASAP.

20 January, 2017

19:00

Came to the observatory.. Roshaan Bukhari (Secretary LAST)  joined me.

We talked about LAST plans for LSM, Astronomy, Astrometry, near future workshops and more.

I tried some new cables to connect and work for me but that did not work out.

Sky was clear for some time but due to high humidity level, the telescope room was closed. 

 

18 January, 2017

21:00

Two first time visitors came to the observatory tonight. Showed them around. We talked about Astronomy reseacrch, Education and Telescopes. It was all great!

13 January, 2017

18:30

A couple of visitors came to the observatory this evening. We saw Venus, Mars, Uranus and Neptune.. along with Orion Nebula and some stars. A lot of Astronomy and Physics was also discussed.

12 January, 2017.

20:00

Some guests were here and we saw a few objects from the C14 and SBIG setup. The moon was destroying the sky but we will managed to observe some. 

23:00

Acquired some data of an asteriod for more than an hour.  

02:00

Tried to find a quasar galaxy but was not successful, probably because it was in low horizon. Also humidity increaed to 89%, resulting in the fog covering the optics and i was forced to close the observatory. 

11 January, 2017

22:00

I came a bit late to the observatory but managed to find Vesta at the right position in The SkyX Pro with a fone guidestar at 359 degrees angle.

Moon is very bright and there is nothing else for me to do but to image bright asteriods.

01:00

Kept imaging Vesta for a long time. The data js unprocessed yet. 

01:30

Observatory is closed. 

25 December, 2016

Came for a short time at the observatory this evening.. a very passionate young pre-engineering student Yausoob Latif wanted to meet me and see the observatory. 

I should him LAST's meeting room and the stiff it has there. Detailed him about the activites we do here every month.

Then we went to the roof, where he saw the three rooms and the telescopes they have. He was very very excited to see all this stuff. We also saw Venus and Mars and some bright stars.

Yausoob wans to learn everything about Astronomy.. I will be happy to help in anyway possible. He will be coming to our monthly meetings from now on. 

15 December, 2016

I have been very busy and could not find time to come to the observatory lately. Yesterday, i came in the evening and saw the damage a mouse has done! Wires were cut and a small amount of Polystyrene sheet were damaged. Found the mouse eventually and got rid of it in about 2 hours! I should keep coming here.

4 Decmeber, 2016

Had visitors at the observatory in the evening. A young and a very enthusiastic child, who is just 5, saw the planets and the moon with the 127 refractor telescope. 

Main observatory was closed. 

3 December, 2016

17:30

In the early evening, two students doing A levels, came to the observatory and i showed them the observatory. Thy are very interested in learning how the telescopes at the observatory works and what can we do with this equipment.

We used Sky Watcher 10 inch Dobsonian telescope to first see the cresent moon. They became very excited! Then we saw nearby and very brilliant Planet Venus. Mars was next and the last was Albireo, the colorful double star.

We also grabbed some quick pictures from the phone camera. 

2 December, 2016

Lahore Astronomical Society had its monthly meeting at the observatory. Telescopes were closed.

30 November, 2016

19:30

Sky is partly cloudy.. and hazy too.

20:45

I have installed the newly arrived GPS receiver in the mount and it is working all fine!

21:37

The humidity is very high.. 79% right now in the observatory.. not good.

22:11

Sky is getting worst here.. I cannot attempt any deep sky objects tonight. I am going to attempt to collect some astrometric data.

22:16

TCF is not focusing.. the focus point has moved beyond 7000.. beyond the limit of TCF. I now need to change the focus manually and readjust few more things.

22:58

Alright.. i have loosened the mirror holding of C14 telescope, changed the TCF value from 7000 to 2500, have manually refocused the scope with the help of Bahtinov mask, collimated the optics again, refocused the system with the automated TCF Focuser and now the star is perfectly focused!

23:07

The humidity level in the observatory room has gone upto 94%!! Things are dripping, literally!

I have turned off everything and closed the roof.